OpenStudy (anonymous):

Relationship A has a greater rate than Relationship B. The table represents Relationship B. Which equation could represent Relationship A? Choose all answers that are correct. Hours worked 2 4 5 8 Amount paid ($) 30.40 60.80 76 121.60 A. B. C. D.

OpenStudy (anonymous):

x1 -x = 4 - 2 x1 - x = 2 x2 - x1 = 5-4 x2 - x1 = 1 y1 - y = 60.80 - 30.40 y1 - y = 30.40 y2 - y1 = 76 - 60.80 y2 - y1 = 15.2 slope = y2-y1/x2-x1 m= 15.20/1 using point slope form y - y2 = m(x - x2) y - 5 = 15.20(x - 76) y = 15.2x - 1155.20 +5 y = 15.2x - 1150.2